# FARIS: Fluid-Active-RIS

### v1

@@ III-B Optimizing $\mathbf{S}_{M_{o}}$ for fixed $\mathbf{v}$
To stochastically determine the active ports for given $\mathbf{v}$, each port $i$ is assigned an activation probability $p_{i}\in[0,1]$, forming the probability vector $\mathbf{p}=[p_{1}\cdots p_{M}]^{\mathrm{T}}$. Let $\bm{\zeta}=[\zeta_{1}\cdots\zeta_{M}]^{\mathrm{T}}\in\{0,1\}^{M}$ denote a random binary port-activation vector sampled as
@@ III-B1 Ideal Target Distribution
Let $\mathcal{Z}\subseteq\{0,1\}^{M}$ denote the set of all feasible binary port-activation vectors, and let $\bar{R}_{S}(\bm{\zeta})$ be the objective in (15) corresponding to $\bm{\zeta}\in\mathcal{Z}$. Define the elite threshold $\gamma_{\rho}$ as the $(1-\rho)$-quantile of $\bar{R}_{S}(\bm{\zeta})$ under the current sampling distribution, so that a fraction $\rho$ of samples satisfy $\bar{R}_{S}(\bm{\zeta})\geq\gamma_{\rho}$. The ideal target distribution $\mathbb{P}^{\dagger}(\bm{\zeta})$ that perfectly concentrates on the elite set is then
@@
Intuitively, $\mathbb{P}^{\dagger}$ assigns nonzero probability only to high-performing (elite) activation patterns.
@@ III-B2 Applying to $M_{o}$-constrained CEM with Iterations
Algorithm 2 CEM-Based Selection of $\mathbf{S}_{M_{o}}$ with Expected $M_{o}$-Cardinality for Fixed $\mathbf{v}$
@@
Algorithm 3 Overall AO Framework for $\mathbf{S}_{M_{o}}$ and $\mathbf{v}$
@@ IV-B3 Gaussian Randomization
The Cholesky factorization of $\mathbf{V}_{0}$ requires $\mathcal{O}(M_{o}^{3})$ flops and is performed once [49]. For each randomized candidate, generating ${\mathbf{v}}_{n}$ and evaluating the $S$ quadratic terms $\mathrm{tr}(\mathbf{C}_{s}\mathbf{V}_{n})$ and $\mathrm{tr}(\mathbf{a}_{s}\mathbf{a}_{s}^{*}\mathbf{V}_{n})$ incur $\mathcal{O}(SM_{o}^{2})$ operations. Furthermore, the reduced conic subproblem in $\{w_{s},\xi_{s}\}$ satisfies $n_{\mathrm{var}}=2S$ and $n_{\mathrm{cone}}=n_{\mathrm{lin}}=3S$; hence the arithmetic cost per interior-point iteration is $\mathcal{O}(K_{\mathrm{IP}}^{\prime}S^{3})$ with $K_{\mathrm{IP}}^{\prime}$ setps by (46). Consequently, the overall complexity of the Gaussian randomization stage is
$\mathcal{O}\big(M_{o}^{3}+N_{\mathrm{rand}}(SM_{o}^{2}+K_{\mathrm{IP}}^{\prime}S^{3})\big),$ (48)
and the element-wise magnitude projection requires $\mathcal{O}(M_{o})$, which is negligible compared to the dominant terms.
@@ V Simulation Results
In the simulations, we assess the effectiveness of the proposed AO frameworks for maximizing $\bar{R}$ through Monte-Carlo evaluations. The FARIS setup, as shown in Fig. 2, considers a deployment of the BS, FARIS, and MU, with distances $l_{f}=3~\mathrm{m}$ and $l_{u}=15~\mathrm{m}$. Since $l_{f}=3~\mathrm{m}$ is smaller than $D^{\mathrm{RB}}=6~\mathrm{m}$, the LoS condition for $\tilde{\mathbf{h}}_{f}$ is satisfied [45]. The channels between the ARIS and MU are modeled as flat Rician fading channels, with a path-loss exponent of 2.2 and $K=1$ [51, 52], and the detailed simulation parameters are listed in Table I. Each Monte-Carlo point is computed by averaging over $10^{3}$ independent realizations of signal transmission and reception. For comparison, we also include the performance of (i) a conventional FRIS (passive) architecture with the same number of $M_{o}$ FRIS elements selected from $M$ candidates with the particle-swarm-optimization (PSO)-based approach in [40], and (ii) an ARIS architecture with same $M_{o}$ elements with sequential convex approximation (SCA)-based optimization in [12]. We additionally compare our results with a BFS baseline that exhaustively evaluates every possible port-phase configuration of the FRIS, assuming a $b=2$-bit discrete phase quantization and one-dimensional search over $0,g_{\max}] [$44].
@@
Reflection power budget of FARIS $P_{\max}$ 25 [dBm] [12]
@@
Noise at FARIS and MU $(\sigma_{r}^{2},\sigma_{0}^{2})$ -90 [dBm]
@@
Parameters in CEM $(N_{\mathrm{mc}},\rho,\omega)$ $(5N,0.1,0.7)$
[50, 42]
Fig. 3 depicts the preset position candidates for the FARIS when employing $M_{o}=16,~36$ and $81$ fluid elements, which show the inherent flexibility of the FARIS architecture. For each case, the most advantageous $\mathbf{S}_{M_{o}}^{\star}$ is determined by CEM-based Algorithm 2. The resulting element selections exhibit a consistent tendency to activate elements located near the array boundary rather than those concentrated toward the center. This behavior can be physically explained by considering the effective aperture and spatial diversity of the FARIS. Elements placed farther from the array centroid experience larger angular separation and distinct incidence geometries, thereby offering lower inter-element correlation and contributing more significantly to the expansion of diversity and DoF through optimized $\mathbf{G}$.
@@
Fig. 4 depicts the cumulative distribution function (CDF) of the ergodic-rate gap, together with that of the BFS benchmark, for $M=16$ and $M_{o}\in\{4,9\}$. Notably, the FARIS equipped with the proposed AO framework achieves near-optimal performance, closely tracking the exhaustive BFS curve despite the dramatic reduction in computational burden. The average gap remains as small as 0.53 and 0.75 bps/Hz for $M_{o}=4$ and $9$, respectively, a practically negligible loss considering that BFS requires evaluating the entire discrete port-phase space and optimal amplification control, whereas our method converges rapidly with orders-of-magnitude lower complexity. This highlights the efficiency and scalability of the proposed AO framework, even under stringent configuration constraints.
@@
Fig. 6 presents $\bar{R}$ versus $P$ for FARIS and conventional FRIS/ARIS under $M_{o}=16,~36$ and $81$. As $P$ increases, FARIS with the proposed AO framework and ARIS initially exhibit a steeper rate growth; however, after certain threshold of $P$, the hardware constraint (9) in (13) becomes active, causing both curves to transition into a more moderate scaling with respect to $P$ [12]. Consequently, the optimization gain of FARIS over ARIS also diminishes slightly in this high-power regime, as the ARIS amplification is limited by the same hardware bound. Nevertheless, FARIS still preserves a substantial rate margin over FRIS/ARIS, even with small $M_{o}$, as its fluid positioning and active amplification enable more effective exploitation of the aperture and yield sustained capacity gains over the counterparts; at $P=25$ dBm, the FARIS with $M_{o}=36$ and $81$ yields 0.36 and 0.73 bps/Hz higher rate compared to ARIS architecture. Furthermore, with same $P=25$ dBm, the FARIS configuration with $M_{o}=16$ attains an average rate of approximately 8.02 bps/Hz, whereas the FRIS setup, despite employing a much larger aperture, achieves no more than 3.82 bps/Hz at $M_{o}=81$, meaning that FARIS provides a 4.20 bps/Hz higher rate. This performance gap widens further for $M_{o}=36$, making the rate of FARIS 5.14 bps/Hz higher. Moreover, increasing $M_{o}$ improves $\bar{R}$ for every architecture; however, the marginal gain decreases for larger $M_{o}$ with fixed $M$, reflecting a saturation due to dense spatial sampling shown in Fig. 3.
@@
Fig. 7 depicts the variation of $\bar{R}$ with respect to $M$ (or equivalently $M_{x}$) in the FARIS and conventional FRIS/ARIS configuration. For any fixed M, FARIS equipped with the proposed AO framework consistently achieves a higher rate than FRIS/ARIS, even with less $M_{o}$; when $M_{x}=16$, the proposed FARIS configuration with $M_{o}=16$ provides an additional 3.66 bps/Hz over FRIS with $M_{o}=81$, which increases to 4.33 bps/Hz when $M_{o}=36$, and the gain over ARIS, whose fixed architecture yields a nearly flat rate profile due to the absence of extra DoFs or adaptive optimization, becomes readily apparent in the figure. This again highlights the substantial rate gain enabled by both fluid repositioning and active amplification. As $M$ increases, both FARIS/FRIS architectures benefit from the denser aperture, which enhances spatial sampling and increases the likelihood of selecting more favorable element positions. Despite this common scaling benefit, FARIS maintains a clear performance margin over FRIS, as its fluid port placement together with active amplification enables more effective utilization of the expanded aperture, thereby providing rate enhancements unattainable by the purely passive counterpart.
@@
Fig. 8 illustrates $\bar{R}$ as a function of $W_{x}$ for both FARIS and conventional FRIS. In the initial growth region, all schemes, including the proposed AO-enabled FARIS and the conventional FRIS, exhibit a sharp increase in $\bar{R}$ due to the rapid reduction in spatial correlation and the improved ability to select favorable port positions as the aperture expands. This effect is also considerably pronounced for FARIS, similar to the previous observation, so that even a FARIS design with a smaller element count (e.g., $M_{o}=16$) can surpass the rate of a larger-aperture FRIS configuration (e.g., $M_{o}=36,~81$). As $W_{x}\gtrapprox 5$, which corresponds to spacing of $\frac{\lambda}{2}$, the rate improvement begins to saturate since the additional aperture provides only marginal decorrelation benefits [29]. Notably, in this region, FARIS exploits the enlarged aperture far more effectively than FRIS, thanks to its jointly optimized fluid positioning and active amplification. Therein with $W_{x}=6$, the rate difference between FARIS with $M_{o}=16$ and FRIS when $M_{o}=81$ is 3.43 bps/Hz, which further widens to 3.90 bps/Hz when the number of active FRIS elements is reduced to $M_{o}=36$. Although all schemes eventually show saturation after $W_{x}\gtrapprox 5$, FARIS maintains a consistently larger gain throughout the entire range of $W_{x}$.
@@ VI Conclusion
In this paper, we proposed the FARIS concept, a novel fluid-active-RIS architecture that jointly exploits fluid port repositioning and per-port amplification to substantially enhance the performance of RIS-aided wireless network. We developed a correlated FARIS channel model based on Jakes’ spatial correlation and formulated an ergodic-rate maximization problem under a realistic reflection-power constraint, jointly optimizing active reflection coefficients and port selection. To tackle the resulting mixed-integer non-convex program, we introduced an AO framework that alternates between: i) a lifted active-reflection design solved via SAA and a quadratic-transform-based conic optimization in the lifted matrix variable, and ii) an $M_{o}$-constrained CEM-based port-selection scheme with a closed-form characterization guaranteeing a unique solution and monotonic improvement of the underlying likelihood. Simulation results showed that FARIS with proposed AO framework converges within a modest number of iterations and consistently outperforms a conventional FRIS/ARIS baseline across a wide range of transmit powers, total elements, and apertures, often achieving higher rates even with fewer active elements or smaller apertures. These findings demonstrate that combining fluid repositioning with active amplification provides a powerful additional DoF over FRIS/ARIS designs, positioning FARIS as a promising building block for 6G wireless environments.

### v2

@@ III-B Optimizing $\mathbf{S}_{M_{o}}$ for fixed $\mathbf{v}$
To stochastically determine the active ports for given $\mathbf{v}$, each port $i$ is assigned an selection probability $p_{i}\in[0,1]$, forming the probability vector $\mathbf{p}=[p_{1}\cdots p_{M}]^{\mathrm{T}}$. Let $\bm{\zeta}=[\zeta_{1}\cdots\zeta_{M}]^{\mathrm{T}}\in\{0,1\}^{M}$ denote a random binary port-selection vector sampled as
@@ III-B1 Ideal Target Distribution
Let $\mathcal{Z}\subseteq\{0,1\}^{M}$ denote the set of all feasible binary port-selection vectors, and let $\bar{R}_{S}(\bm{\zeta})$ be the objective in (16) corresponding to $\bm{\zeta}\in\mathcal{Z}$. Define the elite threshold $\gamma_{\rho}$ as the $(1-\rho)$-quantile of $\bar{R}_{S}(\bm{\zeta})$ under the current sampling distribution, so that a fraction $\rho$ of samples satisfy $\bar{R}_{S}(\bm{\zeta})\geq\gamma_{\rho}$. The ideal target distribution $\mathbb{P}^{\dagger}(\bm{\zeta})$ that perfectly concentrates on the elite set is then
@@
Intuitively, $\mathbb{P}^{\dagger}$ assigns nonzero probability only to high-performing (elite) selection patterns.
@@ III-B2 Applying to $M_{o}$-constrained CEM with Iterations
Algorithm 2 CEM-Based Selection of $\mathbf{S}_{M_{o}}$ for Fixed $\mathbf{v}$
@@
Algorithm 3 Overall AO Framework
@@ IV-B3 Gaussian Randomization
The Cholesky factorization of $\mathbf{V}_{0}$ requires $\mathcal{O}(M_{o}^{3})$ flops and is performed once [31]. For each $\mathbf{z}_{n}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M_{o}})$, $\mathbf{v}_{n}=\mathbf{L}\mathbf{z}_{n}$ is generated with complexity $\mathcal{O}(M_{o}^{2})$. Each $\mathbf{v}_{n}$ is then projected onto the feasible set by (28) and (29), dominated by latter one with $\mathcal{O}(M_{o}^{2})$ operations. For each feasible $\mathbf{v}_{n}$, we form $\mathbf{V}_{n}=\mathbf{v}_{n}\mathbf{v}_{n}^{*}$ and evaluate $\tilde{R}_{n}$ by computing $\{\mathrm{tr}(\mathbf{C}_{s}\mathbf{V}_{n}),\mathrm{tr}(\mathbf{a}_{s}\mathbf{a}_{s}^{*}\mathbf{V}_{n})\}_{s=1}^{S}$ in the objective of (22), which requires $\mathcal{O}(SM_{o}^{2})$ operations; hence, the cost per randomized candidate is $\mathcal{O}(SM_{o}^{2})$. Consequently, the overall computational complexity is
$\mathcal{O}\Big(M_{o}^{3}+N_{\mathrm{rand}}SM_{o}^{2}\Big).$ (49)
@@ V Simulation Results
In the simulations, we assess the effectiveness of the proposed AO frameworks for maximizing $\bar{R}$ through Monte-Carlo evaluations. The FARIS setup, as shown in Fig. 3, considers a deployment of the BS, FARIS, and MU, with distances $l_{f}=3~\mathrm{m}$ and $l_{u}=15~\mathrm{m}$. Since $l_{f}=3~\mathrm{m}$ is smaller than $D^{\mathrm{RB}}=6~\mathrm{m}$, the LoS condition for $\tilde{\mathbf{h}}_{f}$ is satisfied [22]. The channels between the ARIS and MU are modeled as flat Rician fading channels, with a path-loss exponent of 2.2 and $K=1$ [48, 38], and the detailed simulation parameters are listed in Table I. Each Monte-Carlo point is computed by averaging over $10^{3}$ independent realizations of signal transmission and reception. For comparison, we also include the performance of (i) a conventional FRIS (passive) architecture with the same number of $M_{o}$ FRIS elements selected from $M$ candidates with the particle-swarm-optimization (PSO)-based approach in [33], and (ii) an ARIS architecture with same $M_{o}$ elements with sequential convex approximation (SCA)-based optimization in [24]. We additionally compare our results with a BFS baseline that exhaustively evaluates every possible port-phase configuration of the FRIS, assuming a $b=2$-bit discrete phase quantization and one-dimensional search over $0,g_{\max}] [$42].
@@
Hardware power consumption $(P_{c},P_{\mathrm{DC}})$ (-10, -5) [dBm] [24]
Total power budget of FARIS $P_{\max,t}$ 25 [dBm] [24]
@@
Noise at FARIS and MU $(\sigma_{r}^{2},\sigma_{0}^{2})$ -90 [dBm] (same)
@@
Parameters in CEM $(N_{\mathrm{mc}},\rho,\omega)$ $(5N,0.1,0.7)$ [32]
Fig. 4 depicts the preset position candidates for the FARIS when employing $M_{o}=16,~36$ and $81$ fluid elements, which show the inherent flexibility of the FARIS architecture. For each case, the most advantageous $\mathbf{S}_{M_{o}}^{\star}$ is determined by CEM-based Algorithm 2. The resulting element selections show a clear tendency to favor ports located near the boundary of the array rather than those concentrated around the center. This phenomenon can be interpreted from the perspective of the effective aperture and spatial diversity provided by FARIS. In particular, elements positioned farther from the array centroid typically experience larger angular separation and more diverse propagation geometries. As a result, they tend to exhibit lower spatial correlation and contribute more effectively to enlarging the achievable diversity and DoF through the optimized configuration of $\mathbf{G}$.
Fig. 5 depicts the cumulative distribution function (CDF) of the ergodic-rate gap, together with that of the BFS benchmark, for $M=16$ and $M_{o}\in\{4,9\}$. Notably, the FARIS equipped with the proposed AO framework achieves near-optimal performance, closely tracking the exhaustive BFS curve despite the dramatic reduction in computational burden. The average gap remains as small as 0.48 and 0.69 bps/Hz for $M_{o}=4$ and $9$, respectively, a practically negligible loss considering that BFS requires evaluating the entire discrete port-phase space and optimal amplification control, whereas our method converges rapidly with orders-of-magnitude lower complexity. This highlights the efficiency and scalability of the proposed AO framework, even under stringent configuration constraints.
@@
Fig. 7 presents $\bar{R}$ versus $P$ for FARIS and conventional FRIS/ARIS under $M_{o}=16,~36$ and $81$. As $P$ increases, FARIS with the proposed AO framework and ARIS initially exhibit a steeper rate growth; however, after certain threshold of $P$, the hardware constraint (10) in (14) becomes active, causing both curves to transition into a more moderate scaling with respect to $P$ [24]. Consequently, the optimization gain of FARIS over ARIS also diminishes slightly in this high-power regime, as the ARIS amplification is limited by the same hardware bound. Nevertheless, FARIS still preserves a substantial rate margin over FRIS/ARIS, even with small $M_{o}$, as its fluid positioning and active amplification enable more effective exploitation of the aperture and yield sustained rate gains over the counterparts; at $P=25$ dBm, the FARIS with $M_{o}=36$ and $81$ yields 0.36 and 0.73 bps/Hz higher rate compared to ARIS architecture. Furthermore, with same $P=25$ dBm, the FARIS configuration with $M_{o}=16$ attains an average rate of approximately 7.97 bps/Hz, whereas the FRIS setup, despite employing a much larger aperture, achieves no more than 3.93 bps/Hz at $M_{o}=81$, meaning that FARIS provides a 4.04 bps/Hz higher rate. This performance gap widens further for $M_{o}=36$, making the rate of FARIS 4.98 bps/Hz higher. Moreover, increasing $M_{o}$ improves $\bar{R}$ for every architecture; however, the marginal gain decreases for larger $M_{o}$ with fixed $M$, reflecting a saturation due to dense spatial sampling shown in Fig. 4.
@@
Fig. 8 depicts the variation of $\bar{R}$ with respect to $M$ (or equivalently $M_{x}$) in the FARIS and conventional FRIS/ARIS configuration. For any fixed M, FARIS equipped with the proposed AO framework consistently achieves a higher rate than FRIS/ARIS, even with less $M_{o}$; when $M_{x}=16$, the proposed FARIS configuration with $M_{o}=16$ provides an additional 3.53 bps/Hz over FRIS with $M_{o}=81$, which increases to 4.20 bps/Hz when $M_{o}=36$, and the gain over ARIS, whose fixed architecture yields a nearly flat rate profile due to the absence of extra DoFs or adaptive optimization, becomes readily apparent in the figure. This again highlights the substantial rate gain enabled by both fluid repositioning and active amplification. As $M$ increases, both FARIS/FRIS architectures benefit from the denser aperture, which enhances spatial sampling and increases the likelihood of selecting more favorable element positions. Despite this common scaling benefit, FARIS maintains a clear performance margin over FRIS, as its fluid port placement together with active amplification enables more effective utilization of the expanded aperture, thereby providing rate enhancements unattainable by the purely passive counterpart.
@@
Fig. 9 illustrates $\bar{R}$ as a function of $W_{x}$ for both FARIS and conventional FRIS. In the initial growth region, all schemes, including the proposed AO-enabled FARIS and the conventional FRIS, exhibit a sharp increase in $\bar{R}$ due to the rapid reduction in spatial correlation and the improved ability to select favorable port positions as the aperture expands. This effect is also considerably pronounced for FARIS, similar to the previous observation, so that even a FARIS design with a smaller element count (e.g., $M_{o}=16$) can surpass the rate of a larger-aperture FRIS configuration (e.g., $M_{o}=36,~81$). As $W_{x}\gtrapprox 5$, which corresponds to spacing of $\frac{\lambda}{2}$, the rate improvement begins to saturate since the additional aperture provides only marginal decorrelation benefits [39]. Notably, in this region, FARIS exploits the enlarged aperture far more effectively than FRIS, thanks to its jointly optimized fluid positioning and active amplification. Therein with $W_{x}=6$, the rate difference between FARIS with $M_{o}=16$ and FRIS when $M_{o}=81$ is 3.27 bps/Hz, which further widens to 3.73 bps/Hz when the number of active FRIS elements is reduced to $M_{o}=36$. Although all schemes eventually show saturation after $W_{x}\gtrapprox 5$, FARIS maintains a consistently larger gain throughout the entire range of $W_{x}$.
Figure 10: Total power consumption versus $M_{o}$ for the proposed FARIS and conventional ARIS and AF relay architectures.
Fig. 10 illustrates the total power consumption versus $M_{o}$ for the proposed FARIS architecture, conventional ARIS, and AF relay. Herein, the hardware power consumption of the $M_{o}$-element AF relay is modeled as [21].
$P_{\mathrm{hard,AF}}=M_{o}(P_{\mathrm{DAC}}+P_{\mathrm{mix}}+P_{\mathrm{filt}})+P_{\mathrm{syn}},$ (53)
where $P_{\mathrm{DAC}},P_{\mathrm{mix}},P_{\mathrm{filt}}$, and $P_{\mathrm{syn}}$ denote the power consumption of the digital-to-analog converter (DAC), mixer, filter, and frequency synthesizer, respectively. According to the practical parameters reported in [21], the combined circuit power $P_{\mathrm{DAC}}+P_{\mathrm{mix}}+P_{\mathrm{filt}}$ is approximately $18.2$ dBm, while $P_{\mathrm{syn}}=17$ dBm. This hardware power is then combined with the communication-related power for AF relays [50, 6]. As shown in the figure, compared with FARIS, the ARIS architecture consistently exhibits lower total power consumption across the considered range of $M_{o}$. This is mainly due to the structural difference between the two architectures. In ARIS, the active elements are directly deployed at the selected $M_{o}$ positions, resulting in a smaller hardware power term $M_{o}(P_{c}+P_{\mathrm{DC}})$ as well as a reduced reflected signal power $P_{\mathrm{RIS}}$. In contrast, FARIS maintains a larger candidate pool with $M$ ports, which introduces an additional control-circuit power $MP_{c}$ in (9). Nevertheless, the power consumption of FARIS remains significantly lower than that of the AF relay. This is because unlike the AF relay nodes that operate as independent transceivers with dedicated radio-frequency (RF) chains and signal forwarding operations [50, 6], whose hardware power in (53) introduces substantially higher total power, FARIS remains a metasurface-based structure in Fig. 2 that directly manipulates the impinging electromagnetic wave over a distributed aperture. Therefore, FARIS provides a favorable intermediate design that achieves notable rate improvements over ARIS while avoiding the excessive power overhead associated with AF relay architectures.
@@ VI Conclusion
In this paper, we proposed the FARIS concept, a novel fluid-active-RIS architecture that jointly exploits fluid port repositioning and per-port amplification to substantially enhance the performance of RIS-aided wireless networks. To capture the practical hardware behavior, we developed a circuit-aware modeling framework and established a correlated FARIS signal model based on Jakes’ spatial correlation. Under a realistic reflection-power constraint, we formulated an ergodic-rate maximization problem that jointly optimizes the active amplification-reflection coefficients and the discrete selection of fluid-active ports. To solve the resulting mixed-integer non-convex problem, we developed an AO-based framework that efficiently alternates between active-reflection optimization and CEM-based port selection, achieving near-optimal performance with significantly reduced computational complexity. Numerical results demonstrate that the proposed FARIS architecture consistently outperforms conventional baselines, achieving higher ergodic rates even with fewer active elements than the benchmark schemes, thereby highlighting the efficiency of the proposed fluid-active design. These results highlight that the synergy between fluid reconfigurability and active amplification introduces a powerful additional spatial DoF, enabling more efficient channel shaping and positioning FARIS as a promising architectural paradigm for 6G.